OpenStudy (anonymous):

Find the standard equation and graph of a parabola that matches the given set of characteristics. vertex (0,0) directix y = -6

OpenStudy (anonymous):

From http://jwilson.coe.uga.edu/emt725/class/sarfaty/emt669/instructionalunit/parabolas/parabolas.html The standard form is (x - h)2 = 4p (y - k), where the focus is (h, k + p) and the directrix is y = k - p.

OpenStudy (anonymous):

Here h and k are zero (vertex at the origin). So p = 6. So x^2 = 24 y^2

OpenStudy (anonymous):

sorry x^2 = 24 y

OpenStudy (anonymous):

Thanks!!!

OpenStudy (anonymous):

Nice to hear from you. yw

OpenStudy (anonymous):

^_^